assume that blood pressure reading are normally distributed with a mean of 125 and a standard deviation of 6.4. if 64 people are randomly selected, find the probability that their mean blood pressure will be less than 127. round to four decimal places.
Answer #1

Solution :

Given that,

mean = = 125

standard deviation = = 6.4

n = 64

= 125

=  / n = 6.4 / 64=0.8

P( <127 ) = P[( - ) / < (127-125) /0.8 ]

= P(z <2.5 )

Using z table  

= 0.9938   

probability= 0.9938